The internal energy of the certain substance is given by the following equation





u = 3.56 pv + 184





where u is given in kJ/kg, p is in kPa and v is in m3





/kg.





A system composed of 3 kg of this substance expands from an initial pressure of 500





kPa and volume of 0.22 m3





to final pressure 100 kPa in a process in which pressure and





volume are related by pv1.2 = Constant. If expansion is quasi-static, find Q, ∆U and W





for the process.

1
Expert's answer
2021-11-29T19:08:25-0500

"u=3.56pv+184 \\

∆u=u_2-u_1 \\

∆u=3.56(p_2V_2-p_1V_1)\\

p_1V_1^{1.2}=p_2V_2^{1.2}\\

V_2=V_1(\frac{p_1}{p_2})^{\frac{1}{1.2}} \\

V_2=0.22\times(\frac{500}{100})^{\frac{1}{1.2}}=0.841 \; m^3 \\

∆U= 3.56(100\times10^3\times 0.841-500\times10^3\times 0.22) \\

∆U=3.56 (100\times10^3(0.841-5 \times 0.22 ) )\\

∆U=-92.204 \; kJ"

Quasi static process

"W=\smallint p dV =\frac{p_1V_1-p_2V_2}{n-1} \\

W=\frac{500\times10^3\times 0.21-100 \times10^3 \times 0.841}{1.2-1}=104.5 \; J \\

Q=∆U+W=-92.204+104.5=12.296 \; kJ

Q=12.3 \; kJ"